OpenStudy (anonymous):

FREE FAN AND MEDAL!!!!!!!!!!!!!!! What is the solution to the equation? 7x + 6 = –22 A. B. x = –4 C. D. x = 4

OpenStudy (anonymous):

A. x=-64/7 C. x=-16/7

OpenStudy (anonymous):

@TheAsker2002 @TheForgottenDogs @iambatman

OpenStudy (anonymous):

@linn99123

OpenStudy (anonymous):

MUST BE RIGHT ANSWERS IF NOT REPORT AND NO FAN NO MEDAL NO TESTEMONIAL AND NO ADVERTISMENT

OpenStudy (anonymous):

but i f by accident its ok

OpenStudy (anonymous):

all right hold on

OpenStudy (anonymous):

so what are you asking for the equation and is it multi choice

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wat is x

OpenStudy (anonymous):

ok looking now

OpenStudy (anonymous):

it would be 7*-4=-28 +6 =-22 so -4

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What is the solution to the equation? 05_05_LQ_3.gif A. z = 50 B. z = 44 C. z = –2 D. z = – 40

OpenStudy (anonymous):

1-z/5=

OpenStudy (anonymous):

=9

OpenStudy (anonymous):

hold on looking

OpenStudy (anonymous):

i don't know this one

OpenStudy (anonymous):

ok